# Implicit U-KAN2.0: Dynamic, Efficient and Interpretable Medical Image Segmentation

### v1

@@ Abstract
Image segmentation is a fundamental task in both image analysis and medical applications. State-of-the-art methods predominantly rely on encoder-decoder architectures with a U-shaped design, commonly referred to as U-Net. Recent advancements integrating transformers and MLPs improve performance but still face key limitations, such as poor interpretability, difficulty handling intrinsic noise, and constrained expressiveness due to discrete layer structures, often lacking a solid theoretical foundation.In this work, we introduce Implicit U-KAN 2.0, a novel U-Net variant that adopts a two-phase encoder-decoder structure. In the SONO phase, we use a second-order neural ordinary differential equation (NODEs), called the SONO block, for a more efficient, expressive, and theoretically grounded modeling approach. In the SONO-MultiKAN phase, we integrate the second-order NODEs and MultiKAN layer as the core computational block to enhance interpretability and representation power. Our contributions are threefold. First, U-KAN 2.0 is an implicit deep neural network incorporating MultiKAN and second order NODEs, improving interpretability and performance while reducing computational costs. Second, we provide a theoretical analysis demonstrating that the approximation ability of the MultiKAN block is independent of the input dimension. Third, we conduct extensive experiments on a variety of 2D and a single 3D dataset, demonstrating that our model consistently outperforms existing segmentation networks.
Keywords:
Image SegmentationU-type NetworksImplicit Learning
@@ 1 Introduction
In this study, we address the limitations of U-KAN by introducing Implicit U-KAN2.0, which builds on implicit neural networks for improved efficiency and interpretability. Our approach introduces two key novelties: 1) the Second Order Neural ODE (SONO) block, which transforms discrete functions into continuous ones while maintaining constant memory cost, and 2) the SONO-MultiKAN block, which integrates SONO with a tokenised MultiKAN layer for enhanced representation power. Compared to U-KAN, our method redesign the architecture by replacing convolutional blocks with SONO, improving discretisation and stability. Furthermore, we introduce a bottleneck module to refine the flow of information between the encoder and decoder, optimising feature retention and improving overall model performance. Instead of additive skip connections, we employ feature concatenation to preserve richer representations. Additionally, unlike KAN-based U-Nets, which lack full GPU compatibility, our model fully optimised for GPU-based training while maintaining constant memory costs and ensuring scalability. Our contributions: I) We introduce a novel implicit deep neural network powered by our SONO block, which enhances the model’s ability to evolve features continuously, improving both accuracy and stability in segmentation tasks. II) We provide theoretical analysis demonstrating that the MultiKAN block’s approximation, with high expressiveness, is independent of input dimensionality. III) Extensive experiments on multiple 2D and 3D datasets confirm that our model outperforms existing segmentation networks.
@@ 2 Methodology
The first key highlight of our Implicit U-KAN2.0 is using Second-Order Neural Ordinary Differential Equations (NODEs) to model the continuous evolution of feature representations, namely SONO block. This dynamic modelling allows for smoother learning trajectories and faster convergence compared to existing methods. The second-order NODEs governing feature evolution is:
$\mathbb{\ddot{x}}(t)=f(\mathbb{x},\mathbb{\dot{x}},t,\theta_{f}),\quad\mathbb{\dot{x}}(t_{0})=g(x_{0},\theta_{g}),\quad\mathbb{x}(t_{0})=x_{0}.$ (1)
where $f$ and $g$ are both neural networks and $\mathbb{x}(t)$ represents the feature vector. The feature was evaluated as a continuous representation and solved by the ode solver in both forward and backward propagation. In addition, the adjoint method was introduced during backpropagation in order to achieve a constant memory cost. By incorporating velocity into the system and extending the solution space to three dimensions, Implicit U-KAN 2.0 achieves faster convergence to optimal feature representations, enhanced stability, and more precise segmentation boundaries. This is particularly important medical imaging where precise boundary delineation is required. To achieve this, we introduce the velocity term $\mathbf{v}(t)=\mathbf{x}^{\prime}(t)$, reformulating the problem into a system of first-order ordinary differential equations (ODEs):$\mathbb{\dot{x}}(t)=\mathbf{v}(t),\hskip 14.22636pt\mathbb{\dot{v}}(t)=f(\mathbf{v}(t),t,\boldsymbol{\theta})$. This transformation expands the phase space to $[\mathbf{x}(t),\mathbf{v}(t)]^{\top}\in\mathbb{R}^{2n}$, which accelerates convergence by enabling trajectory corrections in both position and velocity enabling trajectory corrections in both position and velocity. As a result, convergence is accelerated while maintaining stability. Additionally, the RK4 method is employed to approximate a more stable solution, further improving numerical robustness. This continuous feature evolution ensures smoother transitions in learned representations, which is crucial for medical image segmentation. By capturing gradual changes, it reduces overfitting and makes the network more resilient to noisy, common in medical imaging. This leads to more stable feature extraction and improved segmentation accuracy, even with low-quality images.
@@
Suppose that we define $\text{MultiKAN}(x)=f$ as in (2) and assume $\Psi_{l,i,j}$ is (k+1)-times continuously differentiable. Then, for a given function $G$-grid B-spline functions, there exist exists a constant $C$, dependent on $f$, such that the following holds for all integers $m$ with $0\leq m\leq k$:
$\bigl{\|}\,\text{MultiKAN}(x)-\bigl{(}\Phi_{L-1}^{G}\circ\Phi_{L-2}^{G}\circ\cdots\circ\Phi_{1}^{G}\circ\Phi_{0}^{G}\bigr{)}x\bigr{\|}_{C^{m}}\;\leq\;C\,G^{-k-1+m}.$ (3)
@@
The architecture of Implicit U-KAN2.0 seamlessly integrates its components within a unified encoder-decoder framework. The operation workflow is delineated into two phases: the SONO Phase and the SONO-MultiKAN Phase. Figure 1 shows the overall architecture and the details component of each Block.
[…]
SONO-MultiKAN Phase: In the SONO-MultiKAN phase, the ODE Block is initially applied, followed by the tokenization process to prepare the features for the MultiKAN layer. First, the SONO output $X_{L}$ from the previous two encoders are transformed into a collection of flattened two-dimensional patches, consistent with the methodology described in [6]. Specifically, $X_{L}$ is partitioned into $M$ patches, each of dimensions $K\times K$, such that $M=\frac{H_{L}\times W_{L}}{K^{2}}$. Each patch, denoted by $X_{L}^{i}\in\mathbb{R}^{k^{2}\cdot C_{L}}$, is subsequently projected into a $d$dimensional embedding space through a learnable linear mapping $E\in\mathbb{R}^{(P^{2}\cdot C_{L})\times d}$. Consequently, the initial token embeddings are represented as $Z_{0}=\bigl{[}X_{L}^{1}E;\;X_{L}^{2}E;\;\dots;\;X_{L}^{N}E\bigr{]}$. Next, a three-layer MultiKAN module is employed to further extract features, which is then followed by a depth-wise convolutional layer (De-WiseConv). A residual connection is incorporated by adding the original token embeddings, mathematically expressed as: $Z_{K}=LN(Z_{k-1}+\text{De-WiseConv}(\text{MultiKAN}((\text{ODEBlock}(X_{k-1})))=LN(Z_{k-1}+\text{De-WiseConv}(\text{MultiKAN}(Z_{k-1})$. This formulation ensures that the enhanced features retain essential information from the initial embeddings while integrating the refined representations obtained through the MultiKAN module.
@@ 4 Conclusion
5 Acknowledgements
CWC is supported by the Swiss National Science Foundation (SNSF) under grant number 20HW-1 220785. YC is funded by an AstraZeneca studentship and a Google studentship. CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skodowska-Curie grant agreement No. 777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information and the Alan Turing Institute. AIAR gratefully acknowledges the support from Yau Mathematical Sciences Center, Tsinghua University.

### v2

@@ Abstract
Image segmentation is a fundamental task in both image analysis and medical applications. State-of-the-art methods predominantly rely on encoder-decoder architectures with a U-shaped design, commonly referred to as U-Net. Recent advancements integrating transformers and MLPs improve performance but still face key limitations, such as poor interpretability, difficulty handling intrinsic noise, and constrained expressiveness due to discrete layer structures, often lacking a solid theoretical foundation.In this work, we introduce Implicit U-KAN 2.0, a novel U-Net variant that adopts a two-phase encoder-decoder structure. In the SONO phase, we use a second-order neural ordinary differential equation (NODEs), called the SONO block, for a more efficient, expressive, and theoretically grounded modeling approach. In the SONO-MultiKAN phase, we integrate the second-order NODEs and MultiKAN layer as the core computational block to enhance interpretability and representation power. Our contributions are threefold. First, U-KAN 2.0 is an implicit deep neural network incorporating MultiKAN and second order NODEs, improving interpretability and performance while reducing computational costs. Second, we provide a theoretical analysis demonstrating that the approximation ability of the MultiKAN block is independent of the input dimension. Third, we conduct extensive experiments on a variety of 2D and a single 3D dataset, demonstrating that our model consistently outperforms existing segmentation networks. Project Website: https://math-ml-x.github.io/IUKAN2/
@@ 1 Introduction
In this study, we address the limitations of U-KAN by introducing Implicit U-KAN2.0, which builds on implicit neural networks for improved efficiency and interpretability. Our approach introduces two key novelties: 1) the Second Order Neural ODE (SONO) block, which transforms discrete functions into continuous ones while maintaining constant memory cost, and 2) the SONO-MultiKAN block, which integrates SONO with a tokenised MultiKAN layer for enhanced representation power. Compared to U-KAN, our method redesigns the architecture by replacing convolutional blocks with SONO, improving discretisation and stability. Furthermore, we introduce a bottleneck module to refine the flow of information between the encoder and decoder, optimising feature retention and improving overall model performance. Instead of additive skip connections, we employ feature concatenation to preserve richer representations. Additionally, unlike KAN-based U-Nets, which lack full GPU compatibility, our model fully optimised for GPU-based training while maintaining constant memory costs and ensuring scalability. Our contributions: I) We introduce a novel implicit deep neural network powered by our SONO block, which enhances the model’s ability to evolve features continuously, improving both accuracy and stability in segmentation tasks. II) We provide theoretical analysis demonstrating that the MultiKAN block’s approximation, with high expressiveness, is independent of input dimensionality. III) Extensive experiments on multiple 2D and 3D datasets confirm that our model outperforms existing segmentation networks.
@@ 2 Methodology
[…]
@@
Suppose that we define $\text{MultiKAN}(x)=f$ as in (1) and assume $\Psi_{l,i,j}$ is (k+1)-times continuously differentiable. Then, for a given function $G$-grid B-spline functions, there exist exists a constant $C$, dependent on $f$, such that the following holds for all integers $m$ with $0\leq m\leq k$: $\bigl\|\,\text{MultiKAN}(x)-\bigl(\Phi_{L-1}^{G}\circ\Phi_{L-2}^{G}\circ\cdots\circ\Phi_{1}^{G}\circ\Phi_{0}^{G}\bigr)x\bigr\|_{C^{m}}\;\leq\;C\,G^{-k-1+m}$.
@@
The architecture of Implicit U-KAN2.0 seamlessly integrates its components within a unified encoder-decoder framework. The operation workflow is delineated into two phases: the SONO Phase and the SONO-MultiKAN Phase. Figure 1 shows the overall architecture and the detailed components of each Block.
SONO Phase: The initial input image $X_{0}$ is first processed to establish an initial velocity. Subsequently, the extracted features are propagated through a SONO Block, yielding an output defined by: $X_{L}=\text{Conv}(\text{ODEBlock}(X_{L-1}))$, where $X_{\ell}\in\mathbb{R}^{H_{\ell}\times W_{\ell}\times C_{\ell}}$, where $X_{\ell}\in\mathbb{R}^{H_{\ell}\times W_{\ell}\times C_{\ell}}$ denotes the feature map at the $l$-th layer. Within the SONO Block, we initially employ a function to model the second-order derivative. The resulting output is then fed into an ODE Block, thereby facilitating the approximation of continuous second-order NODEs. Subsequently, we enhance the number of features and perform downsampling via a convolutional layer. In the forward pass, the ODE block is formulated as $\text{ODEBlock}(v_{t})=\text{ODESolve}(\mathbf{v}(t_{0}),f,t_{0},t_{u},\theta)$ and the loss function is defined by $L(\textbf{v}(t_{u}))=L\left(\textbf{v}(t_{0})+\int_{t_{0}}^{t_{u}}f(\textbf{v}(t),t,\theta)\,dt\right)=L\left(\text{ODESolve}(\textbf{v}(t_{0}),f,t_{0},t_{u},\theta)\right)$.
SONO-MultiKAN Phase: In the SONO-MultiKAN phase, the ODE Block is initially applied, followed by the tokenization process to prepare the features for the MultiKAN layer. MultiKAN utilizes learnable activation functions to enhance interpretability. First, the SONO output $X_{L}$ from the previous two encoders are transformed into a collection of flattened two-dimensional patches, consistent with the methodology described in [6]. Specifically, $X_{L}$ is partitioned into $M$ patches, each of dimensions $K\times K$, such that $M=\frac{H_{L}\times W_{L}}{K^{2}}$. Each patch, denoted by $X_{L}^{i}\in\mathbb{R}^{k^{2}\cdot C_{L}}$, is subsequently projected into a $d$ -dimensional embedding space through a learnable linear mapping $E\in\mathbb{R}^{(P^{2}\cdot C_{L})\times d}$. Consequently, the initial token embeddings are represented as $Z_{0}=\bigl[X_{L}^{1}E;\;X_{L}^{2}E;\;\dots;\;X_{L}^{N}E\bigr]$. These embeddings are then processed by a MultiKAN module with interleaved multiplicative layers to capture higher-order, non-linear feature interactions. Next, a three-layer MultiKAN module is employed to further extract features, which is then followed by a depth-wise convolutional layer (De-WiseConv). A residual connection is incorporated by adding the original token embeddings, mathematically expressed as: $Z_{K}=LN(Z_{k-1}+\text{De-WiseConv}(\text{MultiKAN}((\text{ODEBlock}(X_{k-1})))=LN(Z_{k-1}+\text{De-WiseConv}(\text{MultiKAN}(Z_{k-1})$. The MultiKAN block provides structural transparency through tokenised basis functions with explicit mathematical roles, unlike saliency maps, which offer approximate explanations for black-box models.
@@ 4 Conclusion
4.0.1 Acknowledgments
CWC and JAMZ are supported by the Swiss National Science Foundation under grant number 20HW-1 220785. YC is funded by an AstraZeneca studentship and a Google studentship. CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC grants EP/S026045/1, EP/T003553/1, EP/N014588/1, the Wellcome Innovator Award RG98755 and the Alan Turing Institute. AIAR gratefully acknowledges the support of the Yau Mathematical Sciences Center, Tsinghua University. This work is also supported by the Tsinghua University Dushi Program.
4.0.2 Disclosure of Interests
The authors declare that they have no competing interests.